5. Based on the 2000 Census, the proportion of the California population aged 15 years old or older who are married is p = 0.524. Suppose n = 1000 persons are to be sampled from this population and the sample proportion of married persons ( ) is to be calculated. pˆa) What is the mean of the sampling distribution of ? pˆb) What is the standard deviation of the sampling distribution of ?pˆc) What is the approximate probability that less than 50% of the people in the sample are married?
1
Expert's answer
2019-10-31T09:54:46-0400

Let "X=" the sample proportion of items: "X\sim B(n, p)" X\sim B(n, p).

The mean and variance for the binomial variate is


"mean=\mu=E(X)=np, Var(X)=\sigma^2=np(1-p)"

Given that "n=1000, p=0.524". Then


"mean=\mu=E(X)=np=1000(0.524)=524>10""Var(X)=\sigma^2=np(1-p)=1000(0.524)(1-0.524)=249.424"

"n(1-p)=1000(1-0.524)=476>10"


We can use the normal model


"N\big(p, \sqrt{{p(1-p) \over n}}\big)=N\big(0.524, \sqrt{{0.524(1-0.524) \over 1000}}\big)"

to approximate the sampling distribution of "\hat{p}"

a) What is the mean of the sampling distribution of "\hat{p}"


"mean=0.524"

b) What is the standard deviation of the sampling distribution of "\hat{p}"


"\sigma_{\hat{p}}=\sqrt{{0.524(1-0.524) \over 1000}}\approx0.0158"

c) What is the approximate probability that less than 50% of the people in the sample are married?


"P(\hat{p}<0.5)=P(z<{0.5-0.524 \over 0.0158})\approx P(Z<-1.5190)\approx0.0644"